A rectangular field measures 30 m by 40 m. What is the length of the diagonal, x, from one corner to the opposite corner?
Answer by mananth(16946)

A rectangular field measures 30 m by 40 m. What is the length of the diagonal, x, from one corner to the opposite corner?
the diagonal forms the hypotenuse with the two sides
x^2 = 30^2 _40^2
x^2= 900 +1600
x^2= 2500
x= sqrt 2500
= 25m